OpenStudy (anonymous):

how do I find the domain and range of a circle? I got (x-3)^2+(y+6)^2=25 for the equation PLEASE HELP FINAL EXAM TOMORROW

OpenStudy (dtan5457):

The formula for the domain of a circle is {x| a-r < x < a+r} SO now just plug in a=-3 r=5 -3-5<x<2 -8<x<2 for domain The range will be {y| b-r < y < b+r} b=6 r=5 1<y<11